# MotionCLR: Motion Generation and Training-free Editing via Understanding Attention Mechanisms

### v1

@@ 1 Introduction
Recently, text-driven human motion generation (Ahn et al., 2018; Petrovich et al., 2022; Tevet et al., 2022b; Lu et al., 2024; Guo et al., 2024a; Hong et al., 2022; Wang et al., 2022, 2024) has attracted significant attention in the animation community for its great potential to benefit versatile downstream applications, such as games and embodied intelligence. As the generated motion quality in one inference might be unsatisfactory, interactive motion editing is valued as a crucial task in the community. To provide more interactive editing capabilities, previous works have tried to introduce some human-defined constraints into the editing framework, such as introducing a pre-defined trajectory for a controllable generation (Xie et al., 2024a; Dai et al., 2024; Shafir et al., 2024) or key-frames for motion in-betweening (Chen et al., 2023a; Tang et al., 2022; Harvey et al., 2020).
@@
Despite such progress, the constraints introduced in these works are mainly in-modality (motion) constraints, which require laborious efforts in the real animation creation pipeline. Such interaction fashions strongly restrict the involving humans in the loop of creation. In this work, we aim to explore a natural editing fashion of introducing out-of-modality signals, such as editing texts. For example, when generating a motion with the prompt “a man jumps.”, we can control the height or times of the “jump” action via adjusting the importance weight of the word “jump”. Alternatively, we can also in-place replace the word “jump” into other actions specified by users. Moreover, in this work, we would like to equip the motion generation model with such abilities without re-training.
@@
However, the key limitation of existing motion generation models is that the modeling of previous generative methods lacks explicit word-level text-motion correspondence. This fine-grained cross-modality modeling not only plays a crucial role in text-motion alignment, but also makes it easier for fine-grained editing. To show the problem, we revisit previous transformer-based motion generation models (Tevet et al., 2022b; Zhang et al., 2024b, 2023b; Chen et al., 2023b). The transformer-encoder-like methods (Tevet et al., 2022b; Zhou et al., 2024) treat the textual input as one special embedding before the motion sequence. However, text embeddings and motion embeddings imply substantially different semantics, indicating unclear correspondence between texts and motions. Besides, this fashion over-compresses a sentence into one embedding, which compromises the fine-grained correspondence between each word and each motion frame. Although there are some methods (Zhang et al., 2024b, 2023b) to perform texts and motion interactions via linear cross-attention, they fuse the diffusion timestep embeddings with textual features together in the forward process. This operation undermines the structural text representations and weakens the input textual conditions. Through these observations, we argue that the fine-grained text-motion correspondence in these two motion diffusion fashions is not well considered. Therefore, it is urgent to build a model with good explainability and clear modeling of fine-grained text-motion correspondence.
@@
To resolve these issues, in this work, we propose a motion diffusion model, namely MotionCLR, with a CLeaR modeling of the motion generation process and fine-grained text-motion correspondence. The main component of MotionCLR is a CLR block, which is composed of a convolution layer, a self-attention layer, a cross-attention layer, and an FFN layer. In this basic block, the cross-attention layer is used to encode the text conditions for each word. More specifically, the cross-attention operation between each word and each motion frame models the text-motion correspondence explicitly. Meanwhile, the timestep injection of the diffusion process and the text encoding are modeled separately. Besides, the self-attention layer in this block is designed for modeling the interaction between different motion frames and the FFN layer is a common design for channel mixing.
@@
  • We propose an attention-based motion diffusion model, namely MotionCLR, with clear modeling of the text-aligned motion generation process. MotionCLR achieves comparable generation performance with state-of-the-art methods.
@@
  • Thanks to these observations, we propose a series of interactive motion editing downstream tasks (see Fig. 1) via manipulating attention layer calculations. We additionally explore the potential of our method to perform grounded motion generation when facing failure cases.
@@ 2 Related Work and Contribution
Text-driven human motion generation (Plappert et al., 2018; Ahn et al., 2018; Lin and Amer, 2018; Ahuja and Morency, 2019; Bhattacharya et al., 2021; Tevet et al., 2022a; Petrovich et al., 2022; Hong et al., 2022; Guo et al., 2022b; Zhang et al., 2024b; Athanasiou et al., 2022; Tevet et al., 2022b; Wang et al., 2022; Chen et al., 2023b; Dabral et al., 2023; Yuan et al., 2023; Zhang et al., 2023a; Shafir et al., 2024; Zhang et al., 2023b; Karunratanakul et al., 2023; Jiang et al., 2024; Zhang et al., 2024e; Xiao et al., 2024; Xie et al., 2024a; Lu et al., 2024; Wan et al., 2024; Guo et al., 2024a; Liu et al., 2024; Han et al., 2024; Xie et al., 2024b; Zhou et al., 2024; Petrovich et al., 2024; Barquero et al., 2024; Wang et al., 2024; Huang et al., 2024; Zhang et al., 2024a) uses textual descriptions as input to synthesize human motions. One of the main generative fashions is a kind of GPT-like (Zhang et al., 2023a; Lu et al., 2024; Guo et al., 2024a; Jiang et al., 2024) motion generation method, which compresses the text input into one conditional embedding and predicts motion in an auto-regressive fashion. Besides, the diffusion-based method (Tevet et al., 2022b; Zhang et al., 2024b, 2023b; Zhou et al., 2024; Chen et al., 2023b; Dai et al., 2024) is another generative fashion in motion generation. Note that most work with this fashion also utilizes transformers (Vaswani et al., 2017) as the basic network architecture. Although these previous attempts have achieved significant progress in the past years, the technical design of the explainability of the attention mechanism is still not well considered. We hope this work will provide a new understanding of these details.
@@
Motion editing aims to edit a motion satisfying human demand. Previous works (Dai et al., 2024; Dabral et al., 2023; Kim et al., 2023) attempt to edit a motion in a controlling fashion, like motion inbetweening and joint controlling. There are some other methods (Raab et al., 2023; Aberman et al., 2020b; Jang et al., 2022) trying to control the style of a motion. However, these works are either designed for a specific task or cannot edit fine-grained motion semantics, such as the height or times of a “jump” motion. Raab et al. (2024a) perform motion following via replacing the queries in the self-attention. Goel et al. (2024) propose to edit a motion with an instruction. However, the fine-grained text-motion correspondence in the cross-attention still lacks an in-depth understanding. There are also some methods designed for motion generation (Li et al., 2002) or editing (Lee and Shin, 1999; Holden et al., 2016; Athanasiou et al., 2024), which are limited to adapt to diverse downstream tasks. Compared to motion editing, the field of diffusion-based image editing has been largely explored. Previous studies have achieved exceptional realism and diversity in image editing (Hertz et al., 2023; Han et al., 2023; Parmar et al., 2023; Cao et al., 2023; Tumanyan et al., 2023; Zhang et al., 2023c; Mou et al., 2024; Ju et al., 2024) by manipulating attentions maps. However, relevant interactive editing techniques and observations are still unexplored in the human animation community.
@@
Our key insights and contribution over previous attention-based motion diffusion models (Tevet et al., 2022b; Zhang et al., 2024b, 2023b; Zhou et al., 2024; Chen et al., 2023b; Dai et al., 2024) lie in the clear explainability of the self-attention and cross-attention mechanisms in diffusion-based motion generation models. The cross-attention module in our method models the text-motion correspondence at the word level explicitly. Besides, the self-attention mechanism models the motion coherence between frames. Therefore, we can easily clarify what roles self-attention and cross-attention mechanisms play in this framework, respectively. To the best of our knowledge, it is the first time in the human animation community to clarify these mechanisms in one system and explore how to perform training-free motion editing involving humans in the loop.
@@ 3.1 How Does MotionCLR Model Fine-grained Cross-modal Correspondence?
[…]
@@
  • Cross-attention module plays a crucial role in learning text-motion correspondence in the CLR block. It takes word-level textual embeddings of a sentence for cross-modality interaction, aiming to obtain fine-grained text-motion correspondence at the word level. Specifically, the attention map explicitly models the relationship between each frame and each word, enabling more fine-grained cross-modality controlling (Detailed comparison with previous methods in Sec. C.3).
@@
In summary, in the basic CLR block, we model interactions between frames and cross-modal correspondence, separately and explicitly. More detailed comparisons with previous work are in Sec. C.3. We analyze both self-attention and cross-attention of MotionCLR in following sections.
3.2 Preliminaries of Attention Mechanism
@@
where $\mathbf{Q}\in\mathbb{R}^{N_{1}\times d}$, $\mathbf{K},\mathbf{V}\in\mathbb{R}^{N_{2}\times d}$. Here, $d$ is the embedding dimension of the text or one-frame motion. In the following section, we take $t=0,1,\cdots,T$ as diffusion timesteps, and $f=1,2,\cdots,F$ as the frame number of motion embeddings $\mathbf{X}\in\texttt{R}^{F\times d}$. For convenience, we name $\mathbf{S}=\mathbf{Q}\mathbf{K}^{\top}$ as the similarity matrix and $\texttt{softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d})$ as the attention map in the following sections.
Figure 3: Empirical study of attention mechanisms. We use “a person jumps.” as an example. (a) Key frames of generated motion. (b) The root trajectory along the $Y$-axis (vertical). The character jumps on $\sim 15-40$f, $\sim 60-80$f, and $\sim 125-145$f, respectively. (c) The cross-attention between timesteps and words. The “jump” word is highly activated aligning with the “jump” action. (d) The self-attention map visualization. It is obvious that the character jumps three times. Different jumps share similar local motion patterns.
3.3 Mathematical Preliminaries of Attention Mechanism in MotionCLR
@@
The self-attention mechanism measures the motion similarity of all frames and aims to select the most similar frames in motion features at each place (Detailed diagram in Appendix C).
@@
where $\mathbf{C}\in\mathtt{R}^{L\times d}$ is the textual embeddings of $L$ word tokens, $\mathbf{Q}\in\mathbb{R}^{F\times d}$, $\mathbf{K},\mathbf{V}\in\mathbb{R}^{L\times d}$. Note that $\mathbf{W}\underline{{}_{\mathbf{\star}}}$ in Eq. 2 and Eq. 3 are not the same parameters, but are used for convenience. As shown in Eq. 3, $\mathbf{K}$ and $\mathbf{V}$ are both the transformed text features. Recalling Eq. 1, the matrix multiplication operation between $\mathbf{Q}$ and $\mathbf{K}$ measures the similarity ($\mathbf{S}=\mathbf{QK}^{\top}$) between motion frames and words in a sentence. Similar to that in self-attention, the $\mathtt{softmax}(\cdot)$ operation works as a “soft” $\mathtt{max}(\cdot)$ function to select which transformed word embedding in $\mathbf{V}$ should be selected at each frame. This operation models the motion-text correspondence explicitly. Therefore, we have the second remark.
@@ 3.4 Empirical Evidence on Understanding Attention Mechanisms
To obtain a deeper understanding of the attention mechanism and verify the mathematical analysis of attention mechanisms, we provide some empirical studies on some cases. Due to the page limits, we leave more visualization results for empirical evidence in Appendix D.
As shown in Fig. 3, we take the sentence “a person jumps.” as an example. Besides the keyframe visualization (Fig. 3a), we also visualize the root trajectory along the $Y$-axis (vertical height, in Fig. 3b). As can be seen in Fig. 3, the character jumps at $\sim 15-40$f, $\sim 60-80$f, and $\sim 125-145$f, respectively. Note that, as shown in Fig. 3c, the word “jump” is significantly activated aligning with the “jump” action in the self-attention map. This not only verifies the soundness of the fine-grained text-motion correspondence modeling in MotionCLR, but also meets the theatrical analysis of motion-text ($\mathbf{Q}$-$\mathbf{K}$) similarity. This motivates us to manipulate the attention map to control when the action will be executed. The details will be introduced in Sec. 4.
We also visualize the self-attention map in Fig. 3d. As analyzed in Sec. 3.3, the self-attention map evaluates the similarity between frames. As can be seen in Fig. 3d, the attention map highlights nine areas with similar motion patterns, indicating three jumping actions in total. Besides the time areas that the “jmup” word is activated are aligned with the jumping actions. The highlighted areas in the self-attention map are of line shapes, indicating the taking-off, in-the-air, and landing actions of a jump with different detailed movement patterns.
4 Versatile Applications via Attention Manipulations
(a) Motion (de-)emphasizing via in/de-creasing cross-attention value.
(b) In-place motion replacement via replacing self-attention map.
(c) Motion sequence shifting via shifting cross-attention map.
Figure 4: Motion editing via manipulating attention maps.

### v2

# Pay Attention and Move Better: Harnessing Attention for Interactive Motion Generation and Training-free Editing

@@ 1 Introduction
Recently, text-driven human motion generation (Ahn et al., 2018; Petrovich et al., 2022; Tevet et al., 2022b; Lu et al., 2024; Guo et al., 2024a; Hong et al., 2022; Wang et al., 2022, 2024) has attracted significant attention in the animation community for its great potential to benefit versatile downstream applications, such as games and embodied intelligence. As the generated motion quality in one inference might be misaligned with the users’ intents, interactive motion editing plays a crucial role in the community by introducing humans into the loop of human-machine interaction.
@@
To this end, some attempts have been made to edit or control the generated motion by specifying sparse motion signals, like 3D joint trajectories (Xie et al., 2024; Dai et al., 2024; Shafir et al., 2024) or motion clips (Tang et al., 2022; Harvey et al., 2020). Despite such progress, the constraints introduced in these works are mainly in-modality (motion) constraints, which require laborious efforts in the real animation creation pipeline. Such interaction fashions strongly restrict the involving humans in the loop of creation. In this work, we aim to develop a more user-friendly editing fashion of introducing out-of-modality signals, such as editing texts. For example, when generating a motion with the prompt “a man jumps.”, we can control the height or times of the “jump” action via adjusting the importance weight of the word “jump”. Alternatively, we can also in-place replace the word “jump” with other actions specified by users. The retrieval-based data annotation process of previous work (Athanasiou et al., 2024) not only restricts the semantic editing applications, but also requires expendable labor. In contrast, in this work, we would like to equip the motion generation model with such abilities in a training-free style.
@@
However, the key limitation of existing motion generation models is that the modeling of previous generative methods lacks word-level text-motion correspondence. This fine-grained cross-modality modeling not only plays a crucial role in text-motion alignment, but also makes it easier for fine-grained editing. To show the problem, we revisit previous transformer-based motion generation models (Tevet et al., 2022b; Zhang et al., 2024a, 2023b; Chen et al., 2023). The transformer-encoder-like methods (Tevet et al., 2022b; Zhou et al., 2024) treat the textual input as one special embedding before the motion sequence. However, such integrated text embeddings and motion embeddings imply substantially different semantics, indicating unclear correspondence between the semantics of specific words and motions. Besides, this fashion over-compresses a sentence into one embedding, which compromises the fine-grained correspondence between each word and each motion frame, as evidenced in Fig. 2. Although there are some methods (Zhang et al., 2024a, 2023b) to perform texts and motion interactions via linear cross-attention, they fuse the diffusion timestep embeddings with textual features together in the forward process. This operation undermines the structural text representations and weakens the input textual conditions (Sec. Fig. 2). Through these observations, we argue that the fine-grained text-motion correspondence in these two motion diffusion fashions is not well explored.
@@
To resolve these issues, in this work, we propose a motion diffusion model, namely MotionCLR, with a CLeaR [^1] For clarification, the word “Clear” here means good explainability of the model. modeling of the motion generation process and word-motion correspondence. The main component of MotionCLR is a CLR block, which is composed of a convolution layer, a self-attention layer, a cross-attention layer, and an FFN layer. In this basic block, the cross-attention layer is used to encode the text conditions for each word. More specifically, the cross-attention operation between each word and each motion frame models the text-motion correspondence at the word level. Meanwhile, the timestep injection of the diffusion process and the text encoding are modeled separately. Besides, the self-attention layer in this block is designed for modeling the interaction between different frames and the FFN is a common design for channel mixing.
@@
  • We propose an attention-based motion diffusion model, namely MotionCLR, with clear modeling of the text-aligned motion generation process.
@@
  • Thanks to these observations, we propose a series of interactive motion editing tasks (see Fig. 1) via manipulating attention layers. Importantly, both motion generation and editing are performed in one model. We additionally explore the potential of grounded motion generation when facing hallucination (Sec. 6).
  • We evaluate the generation quality of our method and achieve comparable generation performance with state-of-the-art methods. Besides, the training-free editing methods even outperform baselines requiring specific training in some scenarios.
@@ 2.1 Text-driven Human Motion Generation
Previous text-driven human motion generation (Plappert et al., 2018; Ahn et al., 2018; Lin and Amer, 2018; Ahuja and Morency, 2019; Bhattacharya et al., 2021; Tevet et al., 2022a; Petrovich et al., 2022; Hong et al., 2022; Guo et al., 2022b; Zhang et al., 2024a; Athanasiou et al., 2022; Tevet et al., 2022b; Wang et al., 2022; Chen et al., 2023; Dabral et al., 2023; Yuan et al., 2023; Zhang et al., 2023a; Shafir et al., 2024; Zhang et al., 2023b; Karunratanakul et al., 2023; Jiang et al., 2024; Xie et al., 2024; Lu et al., 2024; Wan et al., 2024; Liu et al., 2024; Zhou et al., 2024; Petrovich et al., 2024; Barquero et al., 2024; Wang et al., 2024) uses textual descriptions as input to synthesize human motions. One of the main generative fashions is a kind of GPT-like (Zhang et al., 2023a; Lu et al., 2024; Guo et al., 2024a; Jiang et al., 2024) motion generation method, which compresses the text input into one conditional embedding and predicts motion in an auto-regressive fashion. Besides, the diffusion-based method (Tevet et al., 2022b; Zhang et al., 2024a, 2023b; Zhou et al., 2024; Chen et al., 2023; Dai et al., 2024) is another generative fashion in motion generation. Note that most work with this fashion also utilizes transformers (Vaswani et al., 2017) as the basic network architecture. Although these previous attempts have achieved significant progress in the past years, the technical design of the explainability of the attention mechanism is still not well considered. We hope this work will provide a new understanding of these details. Besides, previous motion generation models also lack the capability of zero-shot motion editing, which is what we would like to explore in this work.
2.2 Human Motion Editing
@@
The human motion editing task aims to edit a motion satisfying human demand. Previous works (Dai et al., 2024; Dabral et al., 2023; Kim et al., 2023) attempt to edit a motion in a controlling fashion, like motion inbetweening and joint controlling. There are some other methods (Raab et al., 2023; Aberman et al., 2020b; Jang et al., 2022) trying to control the style of a motion. However, these works are either designed for a specific task or cannot edit fine-grained motion semantics, such as the height or times of a “jump” motion. Raab et al. (2024a) perform motion following via replacing the queries in the self-attention, which does not consider the semantic manipulations. Goel et al. (2024) propose to edit a motion with an instruction. However, the MEOs pipeline relies on the text-only LLM outputs, which will introduce hallucinations. MotionFix (Athanasiou et al., 2024) proposes to use the language command to edit motions. However, it needs annotations on the editing text, additionally requiring more labor efforts. COMO (Huang et al., 2025) introduces editing motion via Large Language Models (LLMs) as a translator, which lacks the editing grounds of original motion content. The key reason why the existing method cannot achieve training-free motion editing is that the fine-grained text-motion correspondence in the cross-attention still lacks an in-depth understanding. There are also some methods designed for motion generation (Li et al., 2002) or editing (Lee and Shin, 1999; Holden et al., 2016), which are limited to adapt to diverse downstream tasks.
@@
Our key insights and contribution over previous attention-based motion diffusion models (Tevet et al., 2022b; Zhang et al., 2024a, 2023b; Zhou et al., 2024; Chen et al., 2023; Dai et al., 2024) lie in the clear explainability of the self-attention and cross-attention mechanisms in diffusion-based motion generation models. The cross-attention module in our method models the text-motion correspondence at the word level explicitly. Besides, the self-attention mechanism models the motion coherence between frames. Therefore, we can easily clarify what roles self-attention and cross-attention mechanisms play in this framework, respectively. To the best of our knowledge, it is the first time in the human animation community to clarify these mechanisms in one system and explore how to perform training-free motion editing involving humans in the loop.
2.3 Visual Editing for Image Contents
Image editing in diffusion models has been more explored than motion editing. Previous studies have achieved exceptional realism and diversity in image editing (Hertz et al., 2023; Han et al., 2023; Parmar et al., 2023; Cao et al., 2023; Tumanyan et al., 2023; Zhang et al., 2023c; Mou et al., 2024; Ju et al., 2024) by manipulating attention maps. Especially, although Hertz et al. (2023) proposes to introduce cross-attention into image editing, these techniques and self-attention-based motion editing are still under-explored. However, relevant interactive editing techniques and observations are still unexplored in the human animation community. The basic reason for insufficient exploration for human animation is lacking a fine-grained modeling between words and motions.
@@ 3.1 How Does MotionCLR Model Fine-grained Cross-modal Correspondence?
Figure 2: System overview of MotionCLR architecture. (a) The U-Net-like denoising network is with two CLR blocks before down/up-sampling. (b) The basic CLR block includes four layers, separating the timestep injection and the text condition. (c) The key component is the text-motion cross-attention at the word level.
@@
  • Cross-attention module plays a crucial role in learning text-motion correspondence in the CLR block. It takes word-level textual embeddings of a sentence for cross-modality interaction, aiming to obtain fine-grained text-motion correspondence at the word level. Specifically, the attention map models the relationship between each frame and each word, enabling more fine-grained cross-modality controlling.
@@
In summary, in the basic CLR block, we model interactions between frames and word correspondence, separately in cross-attention. We analyze both self-attention and cross-attention of MotionCLR in the following sections, which is useful for subsequent editing tasks.
3.2 Mathematical Preliminaries of Attention Mechanism
@@
where $\mathbf{Q}\in\mathbb{R}^{N_{1}\times d}$, $\mathbf{K},\mathbf{V}\in\mathbb{R}^{N_{2}\times d}$. Here, $d$ is the embedding dimension of the text or one-frame motion. In the following section, we take $t=0,1,\cdots,T$ as diffusion timesteps, and $f=1,2,\cdots,F$ as the frame number of motion embeddings $\mathbf{X}\in\mathbb{R}^{F\times d}$. For convenience, we name $\mathbf{S}=\mathbf{Q}\mathbf{K}^{\top}$ as the similarity matrix and $\mathbf{A}=\texttt{softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d})$ as the attention map in the following sections.
Figure 3: Empirical study of attention mechanisms. We use “a person jumps.” as an example. (A) Keyframes and the root trajectory of generated motion. The character jumps on $\sim 15-40$f, $\sim 60-80$f, and $\sim 125-145$f, respectively. (B) The cross-attention map between timesteps and words. The “jump” word is highly activated aligning with the “jump” action. (d) The self-attention map visualization. It is obvious that the character jumps three times, reflecting nine areas in the self-attention map. Different jumps share similar local motion patterns.
3.3 Self and Cross Attention Mechanisms in MotionCLR
@@
The self-attention mechanism measures the motion similarity of all frames and aims to select the most similar frames in motion features at each place.
@@
where $\mathbf{C}\in\mathbb{R}^{L\times d}$ is the textual embeddings of $L$ word tokens, $\mathbf{Q}\in\mathbb{R}^{F\times d}$, $\mathbf{K},\mathbf{V}\in\mathbb{R}^{L\times d}$. Note that $\mathbf{W}\underline{{}_{\mathbf{\star}}}$ in Eq. 2 and Eq. 3 are not the same parameters, but are used for convenience. As shown in Eq. 3, $\mathbf{K}$ and $\mathbf{V}$ are both the transformed text features. Recalling Eq. 1, the matrix multiplication operation between $\mathbf{Q}$ and $\mathbf{K}$ measures the similarity ($\mathbf{S}=\mathbf{QK}^{\top}$) between motion frames and words in a sentence. Similar to that in self-attention, the $\mathtt{softmax}(\cdot)$ operation works as a “soft” $\mathtt{max}(\cdot)$ function to select which transformed word embedding in $\mathbf{V}$ should be selected at each frame. This operation models the motion-text correspondence explicitly. Therefore, we have the second remark.
@@ 3.4 Empirical Evidence on Understanding Attention Mechanisms
To obtain a deeper understanding of the attention mechanism and verify the mathematical analysis of attention mechanisms, we provide some empirical studies on some cases.
As shown in Fig. 3, we take the sentence “a person jumps.” as an example. We visualize the keyframe in Fig. 3(A), where we also visualize the root trajectory. As can be seen in Fig. 3(A), the character jumps at $\sim 15-40$f, $\sim 60-80$f, and $\sim 125-145$f, respectively. Note that, as shown in Fig. 3(B), the word “jump” is significantly activated aligning with the “jump” action in the cross-attention map. This not only verifies the soundness of the fine-grained text-motion correspondence modeling in MotionCLR, but also meets the theatrical analysis of motion-text ($\mathbf{Q}$-$\mathbf{K}$) similarity. This motivates us to manipulate the attention map to control when the action will be executed. The details will be introduced in Sec. 4. We also visualize the self-attention map in Fig. 3(C). As analyzed in Sec. 3.3, the self-attention map evaluates the similarity between frames. As can be seen in Fig. 3(C), the attention map highlights nine areas with similar motion patterns, indicating three jumping actions in total. Besides the temporal areas that the “jmup” word is activated are aligned with the jumping actions. The highlighted areas in the self-attention map are line areas, not square areas, indicating the taking-off, in-the-air, and landing actions of a jump with different detailed movement patterns. Due to the page limits, we leave more visualization for empirical evidence in Appendix D.
4 Motion Editing Applications via Attention Manipulations
(a) In-place motion replacement via replacing cross-attention map (Sec. 4.1). The batch size of inference examples is two during the inference stage (reference and edited motion respectively). We replace the cross-attention map of the edited motion as the one of the reference motion.
[…]
(b) Motion (de-)emphasis via in/de-creasing cross-attention value. We can emphasize or de-emphasize a word in the prompt by adjusting the weight of the word. $\alpha\textless 0$ and $\alpha\textgreater 0$ mean emphasis and de-emphasis respectively.
(c) Motion sequence shifting via shifting self-attention map. We adjust the sequence order of the motion sequence via shifting the attention map along the temporal axis.
Figure 4: Diagram of motion editing via manipulating attention maps.